OpenStudy (anonymous):

Will give medal After how many months will Rita pay 200$ in simple annual interest on a credit card that charges on her 15% on her 400$ balance?

OpenStudy (anonymous):

is it 15 months

OpenStudy (anonymous):

formula for simple interest is SI=P*R*T/100 where SI is simple interest,P is principal amount u have at start,r is rate and t is time so 200=400*15*t/100 or t=100/30=3 years and 4 months approx

OpenStudy (anonymous):

oo i seee thank you so much

OpenStudy (anonymous):

i calculated assuming the rate given is a per annum basis

OpenStudy (anonymous):

thanks